you have orders to start a heparin infusion on your patient to start a 20units/kg/hour. your patient weighs 87kg. you have a bag of heparin supplied as 100units/ml. how many ml/hour do you set your IV pump?
Answer #1

ANSWER.

* Calculation of amount(ml/hr) set in IV pump.

              ml/hr = Volume(ml) /Time (hr)

        Order = 20 units/kg/hour.

            Patient weight = 87 kg.

    Amount(ml) of drug required.

   Formula = Desired dose /Dose in hand *Quantity.

         Dose required for the patient weight 87 kg = 20 *87

                                                                        = 1740 unit/hour.

          Dose in hand (H) = 100 unit.

               Quantity(Q) = 1 ml.

     As per formula

              Amount (ml) required = D/H *Q

                                                 = 1740/100 *1

                                                 = 17.4 ml.

       ml/hr set in the IV pump = 17.4/1

                                               = 17.4 ml/hr.